Calculate the molality when 100 g of benzene(C 6 H 6 ) is dissolved in 500 g of hexane

(C 6 H 14 ).


1
Expert's answer
2021-02-03T07:07:03-0500

Mass of benzene(solute)= 100g

Mass of hexane(solvent)= 500g = 0.5Kg

Molar mass of benzene= 78g/mol

Mole of benzene= mass/molar mass

Mole of benzene= 100/78= 1.28mol


Molality= moles of solute/1Kg of solvent

Molality= 1.28/0.5

Molality= 2.56 molal